A chocolate chip cookie manufacturing company recorded the number of chocolate chips in a sample of 50 cookies. The mean is 23.83 and the standard deviation is 2.23 Construct an 80​% confidence interval estimate of the standard deviation of the numbers of chocolate chips in all such cookies.
Answer #1

We have given here,                                              
                                              
                                      
Sample standard deviation=2.23                                  
Sample size =50                                      
                                              
Degree of freedom =n-1=   49                                  
Level of significance given =1-0.80=0.2                                  
Chi square critical value for lower tail=36.82                          
                                              
Chi square critical value for upper tail=62.04                          
                                              
We can use here confidence interval formula for population standard deviation  
(n 1)S2 2 χ(a/2 ) (n 1)S2 <σ< 2 Xi-(a/2)

(50 1) (2.23)2 (50 1)(2.23)2 <σ< 62.04 36.82

=(1.98,2.57 )   
                                              
                                              
                                              
                  
                                              
Lower confidence limit for population standard deviation=1.98                  
                                              
Upper confidence limit for Population standard deviation=2.57